OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, -4) and parallel to y = 3x - 6.

OpenStudy (anonymous):

y = 3x - 10 y = 3x - 6 y = 3x - 2 y = 3x - 12

OpenStudy (anonymous):

Since it said find the parallel equation the slope will stay the same. You have to find b(y intercept using the points they gave you.

OpenStudy (anonymous):

(2, -4) -4 = 3(2) + b

OpenStudy (anonymous):

After you find b, you can easily get the equation.